OpenStudy (anonymous):

evaluate the expression cot(15°)

OpenStudy (mathstudent55):

What are you allowed to use? A calculator?

OpenStudy (mathstudent55):

cot A = 1/tan A

OpenStudy (anonymous):

no I cant use a calculator i have to show my work

OpenStudy (mathstudent55):

Can you find tan(30)?

OpenStudy (anonymous):

Do you know the unit circle?

OpenStudy (anonymous):

tan(30)=1/√3

OpenStudy (campbell_st):

do you know about the sum and difference trig functions...?

OpenStudy (mathstudent55):

Now you can use the half angle formula for tan to get tan(15), Then cot(15) = 1/tan(15)

OpenStudy (campbell_st):

as cot = 1/tan so you could take the reciprocal of \[\cot(45 - 30) = \frac{1}{\tan((45 - 30)} = \frac{1 +\tan(45)\tan(30)}{\tan(45)-\tan(30)}\]

OpenStudy (campbell_st):

then use the exact value triangle for tan(45) and tan(30) then substitute

OpenStudy (anonymous):

how do i find those exact values?

OpenStudy (mathstudent55):

tan(A/2) = (1 - cosA)/(sin A) If you use sin(30) and cos(30) in the formula above, you get tan(15) Then do cot(15) = 1/tan(15)

OpenStudy (campbell_st):

well here are the triangles  hope this helps

OpenStudy (campbell_st):

remember tan = opp/adj

OpenStudy (anonymous):

so im solving for tangent?

OpenStudy (campbell_st):

yep... in the left triangle what is the value of tan(45) in the right triangle, what is the value of tan(30) when you have them substitute them into the formula above..